# A Survey on Failure Analysis and Fault Injection in AI Systems

### v1

@@ 5.3. Gap between Failure Analysis and Fault Injection in AI Model
Failure DeepMutation DeepMutation++ MuNN DeepCrime Covered
Data Quality True
@@
Layer and Neuron Quantity Fault True
@@
Misuse Neural Networks True
@@
Differentiated Focus. The distinct FI tools appear to concentrate on disparate facets of AI model failures. For instance, DeepMutation and DeepCrime are adept at handling data quality and preprocessing faults, whereas MuNN is tailored towards layer and neuron quantity faults. This specialization implies that the selection of an FI tool should be contingent upon the specific fault types targeted for analysis.
Coverage Inconsistency.  Table 6 reveals a disparity in coverage, with certain fault types, such as ”Layer and Neuron Quantity Fault” and ”Misuse Activation Function”, being addressed by multiple tools, while others, like ”Inappropriate LR, Epochs, and BS”, are solely within the purview of DeepCrime. This inconsistency may reflect the inherent challenges associated with simulating specific fault types or indicate a relative lack of focus within the FI.
@@ 6. Failure Analysis and Fault Injection for AI Framework
AI framework layer, including TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2019), and Keras (Keras, 2024), acts as a bridge between the AI model layer and the underlying hardware and system infrastructure. Akin to other software systems, these AI frameworks are susceptible to a variety of faults (Tambon et al., 2024; Li et al., 2023). Failures at the AI framework layer can lead to unavailability, incorrect outputs, and poor performance perceived by users. Thus, ensuring the robustness of AI frameworks is crucial for the reliability of AI systems. This section is dedicated to analyzing faults at the AI framework layer.
@@ 6.1. Failure Analysis in AI Framework
Over recent years, a significant volume of research (Zhang et al., 2018; Abadi et al., 2016; Islam et al., 2019; Humbatova et al., 2020; Chen et al., 2023; Tambon et al., 2024; Li et al., 2023; Yang et al., 2022) has been dedicated to the analysis of failures in AI frameworks. Studies on faults in AI failure can be primarily bifurcated into two categories including failures arising from the usage of AI frameworks, and failures stemming from the frameworks’ implementation. We summarize the types of failures in Table 7 that can occur both in the utilization and the implementation of AI frameworks.
@@
Input Format Fault The shape or type of input data mismatches the expected format. (Humbatova et al., 2020; Islam et al., 2019)
@@
API API Usage Fault API is used in a way that does not conform to the correct logic. (Humbatova et al., 2020)
@@
Configuration Framework Config. Fault Incorrect configuration when using a framework. (Zhang et al., 2018; Yang et al., 2022; Quan et al., 2022)
@@
Environment Config. Fault Environmental configuration faults when developing and deploying an AI framework. (Li et al., 2023; Zhang et al., 2020b; Chen et al., 2023)
Performance Memory Management Fault Faults occur when managing memory between heterogeneous devices. (Makkouk et al., 2022; Yang et al., 2022; Quan et al., 2022)
@@
Configuration Fault typically occurs due to incorrect configuration of the framework. One example of this type of fault in TensorFlow is the confusion with computation model. Users may incorrectly constructed TensorFlow computation graphs using control-flow instead of data-flow semantics (Zhang et al., 2018; Yang et al., 2022). Quan et al. (Quan et al., 2022) also analyse the failures in building and initialing JavaScript-based DL systems, such as npm package installation and multi-backend initialization. Another situation of this fault is the misconfiguration of the computing device (e.g., GPU). This type of misconfiguration can include selecting the wrong GPU device, mistakenly using CPU tensors instead of GPU tensors, or improper allocation of resources between CPU and GPU (Humbatova et al., 2020; Yang et al., 2022).
@@
Environment configuration faults mainly encompass the problems during the development and deployment processes of AI framework. Given that AI frameworks typically function in heterogeneous environments, ensuring compatibility with various devices and systems becomes crucial during the development process (Chen et al., 2023; Yang et al., 2022). This can result in failures during the build and compilation process, which hinders the development of AI frameworks. Apart from encountering environment configuration faults during the development process, deploying an AI framework also entails addressing environment faults, such as ”path not found”, ”library not found” and ”permission denied”  (Zhang et al., 2020b). Moreover, deploying the AI framework on various operating systems (e.g., Linux, Windows, Mac, and Docker environments) or utilizing different types of acceleration devices within the framework can also give rise to environment-related faults (Li et al., 2023).
@@
Performance Fault typically does not result in system downtime but can significantly impact the runtime of the system. In the aspect of AI framework, there is a wide range of causes for performance faults which can be quite diverse. One of the causes is memory inefficiencies. Existing AI frameworks such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016) are typically implemented using C/C++ and CUDA, and their memory management is often done manually (Makkouk et al., 2022; Yang et al., 2022). These frameworks need to handle memory exchanges between heterogeneous devices, which can potentially introduce memory inefficiency faults (Quan et al., 2022). Apart from memory management faults, another cause for performance faults is threading inefficiency (Makkouk et al., 2022). Such fault is commonly found in GPU related code. Insufficient parallelism can result in underutilization of device resources, while excessive parallelism can introduce additional overhead(e.g., context switches). Another cause is the trade-off of using different linear algebra libraries/operators. For example, when performing a small matrix operation on a GPU, the computation time may be longer compared to performing the same operation on a CPU (PyTorch, 2024b).
@@
Code Fault primarily refers to logic faults that occur during the implementation of AI framework. One example of code fault in AI framework is a syntax fault, which may occur both in the utilization and the implementation of AI framework. Expect traditional syntax fault occurring in command software system, AI frameworks also face faults related to tensor syntax faults during the implementation (Chen et al., 2023; Tambon et al., 2024). Such faults may occur on account of tensor shape misalignment and operation on tensors across different devices. Apart from common syntax faults, another noteworthy code fault in AI frameworks is the problem in cross-programming-language communication. This kind of fault is particularly common in AI frameworks that utilize multi-programming-language (Li et al., 2023).
@@
Algorithm Fault is related to the defects in algorithm design (Chen et al., 2023; Tambon et al., 2024; Yang et al., 2022). This algorithm fault can be primarily categorized into two aspects including the incorrect implementation logic of an algorithm and the inefficient algorithm implementation. The former aspect mainly pertains to bugs present in the algorithm implementation within the AI framework (Chen et al., 2023; Tambon et al., 2024; Quan et al., 2022). The latter aspect arises from the challenge faced by AI framework developers in keeping up with the latest research and incorporating the most efficient methods for algorithm implementation (Makkouk et al., 2022; Jia et al., 2020).
@@ 6.2. Fault Injection in AI Framework
In recent years, there has been a significant increase in research focusing on FI techniques specifically targeted at AI frameworks. As shown in Table 8, we elaborate on these works that are categorized according to different AI frameworks. These techniques often rely on a process known as ”instrumentation”. This is a method used in fault injection where the system, such as source code or logic gates, is modified to inject faults more accurately or efficiently.
@@
Tensorflow. There are a series of works focus on designing an FI system for TensorFlow as it is one of the most popular frameworks in AI application. TensorFI (Chen et al., 2020c) introduces an interface-level FI approach that focuses on the data-flow graph of TensorFlow. During the inference phase, TensorFI injects both hardware or software faults into TensorFlow operators, corrupting the output of the affected operators. As AI applications developed using TensorFlow 2 do not necessarily depend on data flow graphs, TensorFI2 (Narayanan et al., 2023) utilizes the Keras API to intercept the state of tensors and injects fault to TensorFlow 2. TensorFI2 employs the Keras Model API to modify the layer state or weight matrices that holds the learned model parameters, and utilizes the Keras backend API to intercept the layer computation or activation matrices that holds the output states of the layers. These make TensorFI2 avoid the overhead of graph duplication and inject faults into the model parameters. InjectTF  (Beyer et al., 2020) is another FI framework designed for TensorFlow. InjectTF implements dedicated FI frameworks for both TensorFlow 1 and TensorFlow 2, namely InjectTF1 and InjectTF2. Similar to TensorFI, InjectTF involves the creation of a new data-flow graph.
@@
PyTorch. TorchFI (Goldstein et al., 2020) is a fault inject tool designed for PyTorch, which simulates bit-flip faults that occur in registers or memory by performing single-bit flips on variables or activations within the framework. TorchFI focuses on convolution and fully connected layers and achieves fault injection by modifying the selected nodes within the neural network. Another related work, PyTorchFI (Mahmoud et al., 2020) allows users to introduce neural network perturbations during the execution phase specifically in convolution operations targeting on weights and neurons in DNN. PyTorchFI does not make any modifications to the neural network topology or the source code of PyTorch itself. Instead, it utilizes PyTorch’s hook functions to perturb the values of neurons during the forward propagation process of the computational model. By utilizing hooks to insert faults, PyTorchFI ensures compatibility with future versions of PyTorch and enables the fault code to run at the native speed of PyTorch, resulting in minimal overhead.
@@
Other Frameworks. As a high-level API of TensorFlow, there are studies that explore fault injection based on Keras (Keras, 2024). SNIFF (Breier et al., 2022) utilizes fault injection to achieve reverse engineering of neural networks. In the experimental process, it investigates and the fault injection to Keras. MindSpore (Mindspore-ai, 2024) is a newly developed open source deep learning computing framework by Huawei. MindFI (Zheng et al., 2021) is capable of performing fault injection on MindSpore at the data, software, and hardware levels, following three concepts including ease of use, stable and efficient.
@@
Framework-agnostic. In contrast to prior approaches, enpheeph (Colucci et al., 2022) does not rely on the underlying DNN frameworks. Consequently, it can seamlessly adapt to different DNN frameworks with minimal, or even zero, modifications to the internal code. enpheeph allows for fault injection at various levels, ranging from bit-level and tensor-level to layer-level. It also provides the flexibility to customize the precise number of bit-accurate injections during the execution process. In terms of network compression, enpheeph is the only framework that comprehensive supports for sparse tensors and quantization. This surpasses the functionality offered by most DNN frameworks, as they generally lack direct support for these features. Furthermore, enpheeph has the ability to scale and expand its capabilities across heterogeneous devices.
@@ 6.3. Gap between Failure Analysis and Fault Injection in AI Framework
Upon comparison of the fault types derived from FA, it becomes evident that there is room for enhancement in existing FI tools for AI frameworks. By identifying the gaps between FI tools and the outcomes of FA, engineers can gain a deeper understanding of the limitations of FI tools.
Fault Type Accommodation. The first gap is related to the fault types that the FI tool needs to accommodate and implement. The fault types identified in FA may not always have corresponding injection implementations. Table 9 presents the fault types supported by existing FI tools in AI framework, along with the fault types revealed during FA. Existing FI tools primarily inject faults by modifying specific tensors and constants or by using bit flipping techniques. Therefore, there are relatively fewer implementations of FI specifically targeting faults that are unrelated to model variables, such as ”Performance Fault” and ”Configuration Fault” which are more commonly encountered in chaos engineering practices targeting microservices. So extensive research and collection of prevalent faults in AI frameworks are required for the development of a FI tool so that the tool can better address the needs and requirements of users (Narayanan et al., 2023).
Framework Divergence. The second gap emerges from the divergence between the AI framework targeted by FA and the framework targeted by FI. This divergence often stems from version differences among AI frameworks. For instance, TensorFlow 1 and TensorFlow 2 demonstrate substantial differences in API usage and runtime logic. Failure analysis conducted on TensorFlow 1 may not be directly applicable to FI tools designed for TensorFlow 2. This implies that the TensorFI in Table 9 is only capable of injecting faults into TensorFlow 1, and not TensorFlow 2. This necessitates taking into account the variations among AI frameworks when designing FI tools and selecting a widely applicable method for FI (Narayanan et al., 2023), which might require a reanalysis of failures.
@@ 7. Failure Analysis and Fault Injection for AI Toolkit
AI toolkit layer acts as an intermediate interface between the AI framework and underlying devices (e.g., GPU and NIC), facilitating the AI framework to utilize external functionalities in its implementation. The most commonly used AI toolkit is CUDA (NVIDIA, 2024a), a parallel computing runtime and API developed by NVIDIA. Besides CUDA, there are other GPU-specific toolkits available for AI and high-performance computing (HPC) solution development. Potential failures at this layer can include unavailability or incorrect outputs. This section delves into FA and FI in AI toolkits

### v2

@@ 5.3. Gap between Failure Analysis and Fault Injection in AI Model
Group Failure DeepMutation DeepMutation++ MuNN DeepCrime Covered
Data Data Quality True
@@
Model Hyperparameter Layer and Neuron Quantity Fault True
@@
Model Structure and Algorithm Misuse Neural Networks True
@@
Differentiated Focus. The distinct FI tools appear to focus on disparate facets of AI model failures. For instance, DeepMutation and DeepCrime are adept at handling data quality and preprocessing faults, whereas MuNN is tailored towards layer and neuron quantity faults. This specialization implies that the selection of an FI tool should be contingent upon the specific fault types targeted for analysis.
Coverage Inconsistency.  Table 7 reveals a disparity in coverage, with certain fault types, such as “Layer and Neuron Quantity Fault” and “Misuse Activation Function”, addressed by multiple tools, while others, like “Inappropriate LR, Epochs, and BS”, are solely within the purview of DeepCrime. This inconsistency may reflect the inherent challenges associated with simulating specific fault types or indicate a relative lack of focus within the FI.
@@ 6. Failure Analysis and Fault Injection for AI Framework
AI framework layer, including TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2019), and Keras (Keras, 2024), acts as a bridge between the AI model layer and the underlying hardware and system infrastructure. Akin to other software systems, these AI frameworks are susceptible to a variety of faults (Tambon et al., 2024; Li et al., 2023). Failures at the AI framework layer can lead to unavailability, incorrect output, and poor performance perceived by users. Thus, ensuring the robustness of AI frameworks is crucial for the reliability of AI systems. This section is dedicated to analyzing faults in the AI framework layer.
@@ 6.1. Failure Analysis in AI Framework
Over recent years, a significant volume of research (Zhang et al., 2018; Abadi et al., 2016; Islam et al., 2019; Humbatova et al., 2020; Chen et al., 2023; Tambon et al., 2024; Li et al., 2023; Yang et al., 2022) has been dedicated to the analysis of failures in AI frameworks. Research on AI failure can be primarily bifurcated into two categories, including failures arising from the usage of AI frameworks and failures stemming from the frameworks’ implementation. We summarize the types of failures in Table 8 that can occur in both the utilization and the implementation of AI frameworks.
@@
Input Format Fault The shape or type of input data mismatches the expected format. (Humbatova et al., 2020; Islam et al., 2019; Wang et al., 2022b)
@@
API API Usage Fault API is used in a way that does not conform to the correct logic. (Humbatova et al., 2020; Wang et al., 2022b)
@@
Configuration Framework Config. Fault Incorrect configuration when using a framework. (Zhang et al., 2018; Yang et al., 2022; Quan et al., 2022; Du et al., 2020)
@@
Environment Config. Fault Environmental configuration faults when developing and deploying an AI framework. (Li et al., 2023; Zhang et al., 2020b; Chen et al., 2023; Du et al., 2020)
Performance Memory Management Fault Faults occur when managing memory between heterogeneous devices. (Makkouk et al., 2022; Yang et al., 2022; Quan et al., 2022; Liu et al., 2022)
@@
Configuration Fault typically occurs due to incorrect configuration of the framework. One example of this type of fault in TensorFlow is the confusion with computation model. Users may incorrectly constructed TensorFlow computation graphs using control-flow instead of data-flow semantics (Zhang et al., 2018; Yang et al., 2022). Quan et al. (Quan et al., 2022) also analyze the failures in building and initializing JavaScript-based DL systems, such as npm package installation and multi-backend initialization. Another situation of this fault is the misconfiguration of the computing device (e.g., GPU). This type of misconfiguration can include selecting the wrong GPU device, mistakenly using CPU tensors instead of GPU tensors, or improper allocation of resources between CPU and GPU (Humbatova et al., 2020; Yang et al., 2022).
@@
Environment configuration faults mainly encompass the problems during the development and deployment processes of AI framework. Given that AI frameworks typically function in heterogeneous environments, ensuring compatibility with various devices and systems becomes crucial during the development process (Chen et al., 2023; Yang et al., 2022). This can result in failures during the build and compilation process, which hinders the development of AI frameworks. Apart from encountering environment configuration faults during the development process, deploying an AI framework also entails addressing environment faults, such as “path not found”, “library not found” and “permission denied”  (Zhang et al., 2020b). Moreover, deploying the AI framework on various operating systems (e.g., Linux, Windows, Mac, and Docker environments) or utilizing different types of acceleration devices within the framework can also give rise to environment-related faults (Li et al., 2023).
@@
Performance Fault typically does not result in system downtime but can significantly impact the runtime of the system. In the aspect of AI framework, there is a wide range of causes for performance faults that can be quite diverse. One of the causes is memory inefficiencies. Existing AI frameworks such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016) are typically implemented using C/C++ and CUDA, and their memory management is often done manually (Makkouk et al., 2022; Yang et al., 2022). These frameworks need to handle memory exchanges between heterogeneous devices, which can potentially introduce memory inefficiency faults (Quan et al., 2022). Apart from memory management faults, another cause of performance faults is threading inefficiency (Makkouk et al., 2022). Such a fault is commonly found in GPU related code. Insufficient parallelism can result in underutilization of device resources, while excessive parallelism can introduce additional overhead (e.g., context switches). Another cause is the trade-off of using different linear algebra libraries/operators. For example, when performing a small matrix operation on a GPU, the computation time may be longer compared to performing the same operation on a CPU (PyTorch, 2024b).
@@
Code Fault primarily refers to logic faults that occur during the implementation of the AI framework. An example of a code fault in AI framework is a syntax fault, which may occur in both the utilization and the implementation of AI framework. Expect traditional syntax fault occurring in command software systems, AI frameworks also face faults related to tensor syntax faults during implementation (Chen et al., 2023; Tambon et al., 2024). Such faults may occur on account of tensor shape misalignment and operation on tensors across different devices. Apart from common syntax faults, another noteworthy code fault in AI frameworks is the problem in cross-programming-language communication. This kind of fault is particularly common in AI frameworks that utilize multi-programming-language (Li et al., 2023).
@@
Algorithm Fault is related to defects in algorithm design (Chen et al., 2023; Tambon et al., 2024; Yang et al., 2022). This algorithm fault can be primarily categorized into two aspects, including the incorrect implementation logic of an algorithm and the inefficient algorithm implementation. The former aspect relates mainly to bugs present in the algorithm implementation within the AI framework (Chen et al., 2023; Tambon et al., 2024; Quan et al., 2022). The latter aspect arises from the challenge faced by AI framework developers in keeping up with the latest research and incorporating the most efficient methods for algorithm implementation (Makkouk et al., 2022; Jia et al., 2020).
@@ 6.2. Fault Injection in AI Framework
In recent years, there has been a significant increase in research focused on FI techniques specifically targeted at AI frameworks. As shown in Table 9, we elaborate on these works that are categorized according to different AI frameworks. These techniques often rely on a process known as “instrumentation”. This is a method used in fault injection in which the system, such as the source code or logic gates, is modified to inject faults more accurately or efficiently.
@@
Tensorflow. There are a series of works focus on designing an FI system for TensorFlow as it is one of the most popular frameworks in AI application. TensorFI (Chen et al., 2020c) introduces an interface-level FI approach that focuses on the data-flow graph of TensorFlow. During the inference phase, TensorFI injects hardware or software faults into TensorFlow operators, corrupting the output of the affected operators. As AI applications developed using TensorFlow 2 do not necessarily depend on data flow graphs, TensorFI2 (Narayanan et al., 2023) utilizes the Keras API to intercept the state of tensors and injects fault to TensorFlow 2. TensorFI2 employs the Keras Model API to modify the layer state or weight matrices that hold the learned model parameters, and utilizes the Keras backend API to intercept the layer computation or activation matrices that holds the output states of the layers. These make TensorFI2 avoid the overhead of graph duplication and inject faults into the model parameters. InjectTF  (Beyer et al., 2020) is another FI framework designed for TensorFlow. InjectTF implements dedicated FI frameworks for both TensorFlow 1 and TensorFlow 2, namely InjectTF1 and InjectTF2. Similarly to TensorFI, InjectTF involves the creation of a new data-flow graph.
@@
PyTorch. TorchFI (Goldstein et al., 2020) is a fault injection tool designed for PyTorch, which simulates bit-flip faults that occur in registers or memory by performing single-bit flips on variables or activations within the framework. TorchFI focuses on convolution and fully connected layers and achieves fault injection by modifying the selected nodes within the neural network. Another related work, PyTorchFI (Mahmoud et al., 2020) allows users to introduce neural network perturbations during the execution phase specifically in convolution operations targeting weights and neurons in DNN. PyTorchFI does not make any modifications to the neural network topology or the source code of PyTorch itself. Instead, it utilizes PyTorch’s hook functions to perturb the values of neurons during the forward propagation process of the computational model. By utilizing hooks to insert faults, PyTorchFI ensures compatibility with future versions of PyTorch and enables the fault code to run at the native speed of PyTorch, resulting in minimal overhead.
@@
Other Frameworks. As a high-level API of TensorFlow, there are studies that explore fault injection based on Keras (Keras, 2024). SNIFF (Breier et al., 2022) utilizes fault injection to achieve reverse engineering of neural networks. In the experimental process, it investigates and the fault injection to Keras. MindSpore (Mindspore-ai, 2024) is a newly developed open-source deep learning computing framework by Huawei. MindFI (Zheng et al., 2021) is capable of performing fault injection in MindSpore at the data, software, and hardware levels, following three concepts that include ease of use, stability, and efficiency.
@@
Framework-agnostic. In contrast to prior approaches, enpheeph (Colucci et al., 2022) does not rely on the underlying DNN frameworks. Consequently, it can seamlessly adapt to different DNN frameworks with minimal, or even zero, modifications to the internal code. enpheeph allows for fault injection at various levels, ranging from bit-level and tensor-level to layer-level. It also provides the flexibility to customize the precise number of bit-accurate injections during the execution process. In terms of network compression, enpheeph is the only framework that comprehensively supports sparse tensors and quantization. This surpasses the functionality offered by most DNN frameworks, as they generally lack direct support for these features. Furthermore, enpheeph has the ability to scale and expand its capabilities across heterogeneous devices.
@@ 6.3. Gap between Failure Analysis and Fault Injection in AI Framework
Upon comparison of the fault types derived from FA, it becomes evident that there is room for improvement in existing FI tools for AI frameworks. By identifying the gaps between the FI tools and the outcomes of the FA, engineers can gain a deeper understanding of the limitations of FI tools.
Fault Type Accommodation. The first gap is related to the fault types that the FI tool needs to accommodate and implement. The fault types identified in FA may not always have corresponding injection implementations. Table 10 presents the fault types supported by the existing FI tools in the AI framework, together with the fault types revealed during FA. Existing FI tools primarily inject faults by modifying specific tensors and constants or by using bit flipping techniques. Therefore, there are relatively fewer implementations of FI specifically targeting faults that are unrelated to model variables, such as “Performance Fault” and “Configuration Fault” which are more commonly encountered in chaos engineering practices targeting microservices. Therefore, extensive research and collection of prevalent faults in AI frameworks are required for the development of a FI tool so that the tool can better address the needs and requirements of users (Narayanan et al., 2023).
Framework Divergence. The second gap emerges from the divergence between the AI framework targeted by FA and the framework targeted by FI. This divergence often stems from version differences among AI frameworks. For instance, TensorFlow 1 and TensorFlow 2 demonstrate substantial differences in API usage and runtime logic. Failure analysis conducted on TensorFlow 1 may not be directly applicable to FI tools designed for TensorFlow 2. This implies that the TensorFI in Table 10 is only capable of injecting faults into TensorFlow 1, and not TensorFlow 2. This necessitates taking into account the variations between AI frameworks when designing FI tools and selecting a widely applicable method for FI (Narayanan et al., 2023), which might require a reanalysis of failures.
@@ 7. Failure Analysis and Fault Injection for AI Toolkit
AI toolkit layer acts as an intermediate interface between the AI framework and underlying devices (e.g., GPU and NIC), facilitating the AI framework to utilize external functionalities in its implementation. The most commonly used AI toolkit is CUDA (NVIDIA, 2024a), a parallel computing runtime and API developed by NVIDIA. In addition to CUDA, there are other GPU-specific toolkits available for AI and high-performance computing (HPC) solution development. Potential failures at this layer can include unavailability or incorrect outputs. This section delves into FA and FI in AI toolkits